OpenStudy (anonymous):

One plane can travel 20 miles per hour faster than another. One of them goes 285 miles in same time it takes the other to go 255 miles. What are their speeds? @MrNood

OpenStudy (mrnood):

distance = speed*time 285 = S1*t 255 = S2 *t s1=s2+20

OpenStudy (anonymous):

How would I set it up so I could solve it?

OpenStudy (anonymous):

@paki , help please?

OpenStudy (paki):

the planes are traveling for 1.2 hours because one has travelled 20 miles faster than the other one... so now... 285mil/1.2hrs= 238mph 255mil/1.2hrs= 212mph

OpenStudy (anonymous):

285 = t(s+20) 255 = ts

OpenStudy (anonymous):

How would I set up the question so I could start working out the answer?

OpenStudy (anonymous):

285 = ts+20t ts = 285 - 20t 255 = 285 - 20t 20t = 285 - 255 20t = 30 t = 1.5 Subbing that in into t for the second equation, 255/1.5 = 170, and the other 190.

OpenStudy (anonymous):

Thank you!

OpenStudy (mrnood):

@paki - you can see that your answer is not correct - the question tells you the speeds differ by 20 mph - but in your answer it is 26mph. I think your assertion about the time (1.2. hours)